OpenStudy (anonymous):

write an equation in slope intercept form for (4, -3)

OpenStudy (anonymous):

use the equation y=mx+b

OpenStudy (anonymous):

-3=m(4)+b m=-1/4 b=-3/4 final equation: y=-1/4(x)-3/4